Calculate net value added at factor cost of:
Purchase machinery: 100
Sales:200
Intermediate cost:90
Indirect tax:12
Change in stock:10
Excuse duty:6
Stock of raw material:5
1
Expert's answer
2019-03-25T08:20:06-0400

Net value added = gross value added minus consumption of fixed capital (depreciation)

Gross value added - the difference between the release of goods and services and intermediate cost.


GVA(in basic prices)= 200-90=110

GVA(at market prices)=GVA(in basic prices) + IT+ED=110+12+6=128


NVA= GVA(at market prices) - D=128-100=28